### v5

@@ 6 Related Work
Sparse attention has received sustained research interest since the introduction of the Transformer architecture vaswani2023attentionneed. Early works are primarily motivated by improving the $O(L^{2})$ complexity of attention with sparse factorizations child2019generatinglongsequencessparse, fixed sparse patterns beltagy2020longformerlongdocumenttransformer, or locality-sensitive hashing kitaev2020reformerefficienttransformer. However, such methods require training from scratch with these modifications, leading to such methods falling out of favor as inroads to better model performance were made by simply scaling pretraining tokens and parameter counts. Recent SoTA models have found some success by applying sliding-window attention to a subset of their layers gemmateam2024gemma2improvingopen; openai2025gptoss120bgptoss20bmodel.
@@
In response to the increased cost of LLM training, researchers have developed a variety of training-free KV cache compression methods. StreamingLLM xiao2024efficientstreaminglanguagemodels was originally developed as a method to enable LLMs to generate tokens beyond the maximum sequence length encountered during training, and allows for coherent text generation with a fixed KV cache budget, although accuracy may suffer. Quest tang2024questqueryawaresparsityefficient extends PagedAttention kwon2023efficientmemorymanagementlarge by offloading KV pages with low attention scores based on pre-computed statistics. InfLLM xiao2024infllmtrainingfreelongcontextextrapolation similarly offloads KV cache blocks to host memory and operates with a similar KV cache structure as SnapStream, but looks up Top-$K$ token blocks for every newly decoded token. SAGE-KV wang2025llmsknowdropselfattention and SnapKV li2024snapkvllmknowslooking perform one-time and length-thresholded KV cache compression respectively, but do not distinguish between prefill and decoding. SnapStream builds on training-free compression literature by combining the one-time compression of SnapKV with efficient decoding using a fixed budget from StreamingLLM, fitting each technique into the prefill-decode scheduling of a production LLM server. We leave incorporating pre-computed, offloaded KV blocks/pages in the style of Quest and InfLLM to future work.
@@
However, training-free approaches invariably lead to some accuracy degradation, leading to a resurgence in trainable sparse attention. Cartridges eyuboglu2025cartridgeslightweightgeneralpurposelong propose compressing large corpuses into trainable KV caches with much smaller sequence lengths that may be optionally loaded using cached prefix lookups. Titans behrouz2024titanslearningmemorizetest propose to augment Transformers with a neural memory module that aims to memorize the data seen during training and improve efficiency and accuracy on long context tasks. DeepSeek-V2 introduced Multi-Head Latent Attention (MLA) deepseekai2024deepseekv2strongeconomicalefficient, which introduces shared low-rank projections of the query, key, and value vectors to reduce KV cache memory along the hidden dimension. However, doing so requires increased compute to project these compressed KVs back into the full hidden dimension during decoding. Native Sparse Attention (NSA) yuan2025nativesparseattentionhardwarealigned proposes a similar KV cache structure to SnapStream, with hierarchical compression of past tokens, that is trainable and aims to reduce complexity for both training and inference. DeepSeek Sparse Attention (DSA) deepseekv32exp proposes to finetune an existing dense attention model instead of training from scratch by training a token selection module, which still computes $O(L^{2})$ scores between tokens, but with fewer heads and smaller hidden dimensions than full attention, thus reducing the compute costs of prefill. While SnapStream already accommodates MLA by compressing the smaller, low-rank latent vectors along the sequence dimension instead of the full-rank KVs, admitting the other architectural modifications mentioned above requires additional modifications to fused kernels that we leave to future work.
7 Conclusion
@@
In this paper, we present SnapStream, a training-free KV cache compression method that employs SnapKV compression during prefill and a StreamingLLM rolling KV cache window during decoding. We describe the details of a static graph implementation of this method, as well as the details for a production 16-socket mixed tensor-parallel and data-parallel mapping of DeepSeek-R1-0528, on SN40Ls. This mapping uses SnapStream to reduce memory pressure and increase decoding throughput. We verify that SnapStream introduces minimal accuracy degradation on long sequence tasks for both Llama 3.1 8B Instruct and DeepSeek-R1, as well as additional reasoning and coding tasks for the latter model. Finally, we show how our SnapStream mapping enables us to achieve a $4.3\times$ improvement in decoding throughput, with at most a 5% increase in prefill latency. We hope this work inspires further study of KV cache compression methods in production LLM serving workloads and a variety of hardware accelerators, and leave a closer examination of hyperparameters and mappings of other sparse attention mechanisms to future work.
Appendix A SnapStream Static Graph Pseudo-code
In this section, we show and describe PyTorch-style pseudocode for the operation of a SnapStream KV cache during prefill and decoding. We compare a naïve implementation of SnapStream with our ring buffer instantiation.
A.1 Rolling Recent Window
Listing 1 shows a naïve implementation of the SnapStream KV cache compression process with a rolling window buffer, similar to StreamingLLM. It uses the prefill_snapstream_topk method described in Listing 6 to extract the top-$K$ evicted tokens between the sink and recent tokens. Construction of the compressed KV cache is then a straightforward concatenation in Lines 25-26.
1def prefill_snapstream_naive_compression(key_states, value_states, attn_weights,
2 sink_token_length, topk_length, recent_token_length,
3 snapkv_observation_length, snapkv_kernel_size, prefill_length):
4 ’’
5 key_states: (batch, n_head, max_seq_len, key_head_dim)
6 value_states: (batch, n_head, max_seq_len, value_head_dim)
7 attn_weights: (batch, n_head, max_seq_len, max_seq_len)
8 sink_token_length: int
9 topk_length: int
10 recent_token_length: int
11 snapkv_observation_length: int
12 snapkv_kernel_size: int
13 prefill_length: int
14 ’’
15 sink_token_keys = key_states[:, :, :sink_token_length, :]
16 sink_token_values = value_states[:, :, :sink_token_length, :]
17 lhb = max(sink_token_length, prefill_length - recent_token_length)
18 recent_token_keys = key_states[:, :, lhb:, :]
19 recent_token_values = value_states[:, :, lhb:, :]
20 topk_keys, topk_values = prefill_snapstream_topk(key_states, value_states,
21 attn_weights, sink_token_length, topk_length, recent_token_length,
22 snapkv_observation_length, snapkv_kernel_size, prefill_length)
23 valid_topk = min(max(prefill_length - (sink_token_length + recent_token_length), 0),
24 topk_length)
25 compressed_keys = cat((sink_token_keys, recent_token_keys, topk_keys), dim=2)
26 compressed_values = cat((sink_token_values, recent_token_values, topk_values), dim=2)
27 return compressed_keys, compressed_values, valid_topk
Listing 1: Naïve SnapStream Prefill KV Cache Compression
Listing 2 shows the corresponding naïve implementation of the SnapStream decoding phase KV cache update. In the case where cache_position (sequence length $L$ in previous sections) is less than max_kv_length, the update is a straightforward scatter (Lines 19-22). In the other case (Lines 24-32), we construct a new KV cache by concatenating slices of the sink, Top-$K$, and recent KVs. To accommodate different cache_position values at different batch indices in a static graph, we perform both KV cache updates and select between the two depending on the batch-indexed cache position on Line 34.
[…]
1def decode_snapstream_naive(compressed_keys, compressed_values, decode_keys,
2 decode_values, cache_position, sink_token_length, topk_length,
3 recent_token_length, valid_topk):
4 ’’
5 compressed_keys: (batch, n_head, compressed_max_seq_len, key_head_dim) bf16
6 compressed_values: (batch, n_head, compressed_max_seq_len, value_head_dim) bf16
7 decode_keys: (batch, n_head, 1, key_head_dim) bf16
8 decode_values: (batch, n_head, 1, value_head_dim) bf16
9 cache_position: (batch, 1, 1) int32
10 sink_token_length: int
11 topk_length: int
12 recent_token_length: int
13 valid_topk: int
14 ’’
15 max_kv_length = sink_token_length + recent_token_length + valid_topk
16 # Case 1: L < max_kv_length
17 cache_position_bounded = where(cache_position >= max_kv_length, max_kv_length - 1,
18 cache_position)
19 case1_compressed_keys = scatter(input=compressed_keys, dim=2,
20 index=cache_position_bounded, src=decode_keys)
21 case1_compressed_values = scatter(input=compressed_values, dim=2,
22 index=cache_position_bounded, src=decode_values)
23 # Case 2: L >= max_kv_length
24 sink_keys = compressed_keys[:, :, :sink_token_length, :]
25 sink_values = compressed_keys[:, :, :sink_token_length, :]
26 topk_keys = compressed_keys[:, :, -valid_topk:, :]
27 topk_values = compressed_values[:, :, -valid_topk:, :]
28 new_recent_keys = compressed_keys[:, :, sink_token_length + 1:-topk_length, :]
29 new_recent_values = compressed_values[:, :, sink_token_length + 1:-topk_length, :]
30 case2_compressed_keys = cat([sink_keys, new_recent_keys, decode_keys, topk_keys])
31 case2_compressed_values = cat([sink_values, new_recent_values, decode_values,
32 topk_values])
33 # select between case 1 and case 2
34 selection_condition = cache_position.squeeze(-1, -2) < max_kv_length
35 compressed_keys = where(selection_condition, case1_compressed_keys,
36 case2_compressed_keys)
37 compressed_values = where(selection_condition, case1_compressed_values,
38 case2_compressed_values)
39 return compressed_keys, compressed_values
Listing 2: Naïve SnapStream Decoding KV Cache Update
A.2 Ring Buffer
To avoid these issues, we implement the KV cache as a ring buffer instead, as described in Section 3.2. Listing 3 shows torch-style pseudocode implementing the construction of such a ring buffer during SnapStream prefill.
1def prefill_streamingllm_slicing(key_states, value_states, sink_token_length,
2 recent_token_length, input_length):
3 ’’
4 key_states: (batch, n_head, max_seq_len, key_head_dim)
5 value_states: (batch, n_head, max_seq_len, value_head_dim)
6 sink_token_length: int
7 recent_token_length: int
8 input_length: int
9 ’’
10 sink_k = key_states[:, :, :sink_token_length, :]
11 sink_v = value_states[:, :, :sink_token_length, :]
12 # calculate left hand bound for ring buffer ranges
13 range2_lhb_factor = (input_length - sink_token_length +
14 recent_token_length) // recent_token_length
15 range2_lhb = (range2_lhb_factor * recent_token_length +
16 sink_token_length - recent_token_length)
17 range2_lhb = max(range2_lhb, sink_token_length)
18 range1_lhb = range2_lhb - recent_token_length
19 range1_lhb = max(range1_lhb, sink_token_length)
20 # selection between range 1 and range 2 for ring buffer construction
21 select_offset = input_length - range2_lhb
22 range2_indices = arange(start=range2_lhb,
23 end=range2_lhb + recent_token_length)[None, None, :]
24 range1_indices = arange(start=range1_lhb,
25 end=range1_lhb + recent_token_length)[None, None, :]
26 range2_k = gather(key_states, range2_indices, dim=2)
27 range1_k = gather(key_states, range1_indices, dim=2)
28 range2_v = gather(value_states, range2_indices, dim=2)
29 range1_v = gather(value_states, range1_indices, dim=2)
30 ringbuffer_k = where(arange(recent_token_length)[None, None, :] <= select_offset,
31 range2_k, range1_k)
32 ringbuffer_v = where(arange(recent_token_length)[None, None, :] <= select_offset,
33 range2_v, range1_v)
34 streamingllm_k = cat(sink_k, ringbuffer_k)
35 streamingllm_v = cat(sink_v, ringbuffer_v)
36 return streamingllm_k, streamingllm_v
Listing 3: SnapStream Prefill Ring Buffer Construction
The code in Listing 3 is then used as a subroutine in Listing 4 to construct the ring buffer for the full SnapStream KV Cache during prefill. This greatly simplifies the decoding stage to a single scatter for each of the key and value tensors, as can be seen in Listing 5.
1def prefill_snapstream_compression(key_states, value_states, attn_weights,
2 sink_token_length, topk_length, recent_token_length,
3 snapkv_observation_length, snapkv_kernel_size, prefill_length):
4 ’’
5 key_states: (batch, n_head, max_seq_len, key_head_dim)
6 value_states: (batch, n_head, max_seq_len, value_head_dim)
7 attn_weights: (batch, n_head, max_seq_len, max_seq_len)
8 sink_token_length: int
9 topk_length: int
10 recent_token_length: int
11 snapkv_observation_length: int
12 snapkv_kernel_size: int
13 prefill_length: int
14 ’’
15 streamingllm_keys, streamingllm_values = prefill_streamingllm_slicing(key_states,
16 value_states, sink_token_length, recent_token_length, prefill_length)
17 topk_keys, topk_values = prefill_snapstream_topk(key_states, value_states,
18 attn_weights, sink_token_length, topk_length, recent_token_length,
19 snapkv_observation_length, snapkv_kernel_size, prefill_length)
20 valid_topk = min(max(prefill_length - (sink_token_length + recent_token_length), 0),
21 topk_length)
22 compressed_keys = cat((streamingllm_keys, topk_keys), dim=2)
23 compressed_values = cat((streamingllm_values, topk_values), dim=2)
24 return compressed_keys, compressed_values, valid_topk
Listing 4: SnapStream Prefill KV Cache Compression
1def decode_snapstream_ring(compressed_keys, compressed_values, decode_keys,
2 decode_values, sink_token_length, recent_token_length, decode_length):
3 ’’
4 compressed_keys: (batch, n_head, compressed_max_seq_len, key_head_dim) bf16
5 compressed_values: (batch, n_head, compressed_max_seq_len, value_head_dim) bf16
6 decode_keys: (batch, n_head, 1, key_head_dim) bf16
7 decode_values: (batch, n_head, 1, value_head_dim) bf16
8 sink_token_length: int
9 recent_token_length: int
10 decode_length: int
11 ’’
12 if decode_length < sink_token_length:
13 ring_buffer_index = decode_length
14 else:
15 ring_buffer_index = (decode_length - sink_token_length +
16 recent_token_length) % recent_token_length + sink_token_length
17 compressed_keys = scatter(compressed_keys, index=ring_buffer_index, src=decode_keys)
18 compressed_values = scatter(compressed_values, index=ring_buffer_index,
19 src=decode_values)
20 return compressed_keys, compressed_values
Listing 5: SnapStream Decoding KV Cache Update
1def prefill_snapstream_topk(key_states, value_states, attn_weights, sink_token_length,
2 topk_length, recent_token_length, snapkv_observation_length,
3 snapkv_kernel_size, prefill_length):
4 ’’
5 key_states: (batch, n_head, max_seq_len, key_head_dim)
6 value_states: (batch, n_head, max_seq_len, value_head_dim)
7 attn_weights: (batch, n_head, max_seq_len, max_seq_len)
8 sink_token_length: int
9 topk_length: int
10 recent_token_length: int
11 snapkv_observation_length: int
12 snapkv_kernel_size: int
13 prefill_length: int
14 ’’
15 attn_weights_sum = attn_weights[:, :,
16 prefill_length - snapkv_observation_length:prefill_length,
17 :prefill_length - snapkv_observation_length].sum(dim=2)
18 snapkv_attn_weights = pool2d(attn_weights_sum, kernel_size=snapkv_kernel_size,
19 stride=1, padding=snapkv_kernel_size//2)
20 snapkv_attn_weights = snapkv_attn_weights[:, :,
21 sink_token_length:-recent_token_length]
22 # (batch, n_head, topk_length)
23 topk_indices = topk(snapkv_attn_weights, topk_length, dim=-1)
24 topk_keys = gather(key_states, sink_token_length + topk_indices, dim=2)
25 topk_values = gather(value_states, sink_token_length + topk_indices, dim=2)
26 return topk_keys, topk_values
Listing 6: SnapStream Prefill SnapKV Compression

### v6

@@ VI Related Work
Sparse attention has received sustained research interest since the introduction of the Transformer architecture [28]. Early works are primarily motivated by improving the $O(L^{2})$ complexity of attention with sparse factorizations [7], fixed sparse patterns [5], or locality-sensitive hashing [18]. However, such methods require training from scratch with these modifications, leading to such methods falling out of favor as inroads to better model performance were made by simply scaling pretraining tokens and parameter counts. Recent SoTA models have found some success by applying sliding-window attention to a subset of their layers [26, 23].
@@
In response to the increased cost of LLM training, researchers have developed a variety of training-free KV cache compression methods. StreamingLLM [32] was originally developed as a method to enable LLMs to generate tokens beyond the maximum sequence length encountered during training, and allows for coherent text generation with a fixed KV cache budget, although accuracy may suffer. Quest [27] extends PagedAttention [19] by offloading KV pages with low attention scores based on pre-computed statistics. InfLLM [31] similarly offloads KV cache blocks to host memory and operates with a similar KV cache structure as SnapStream, but looks up Top-$K$ token blocks for every newly decoded token. SAGE-KV [29] and SnapKV [20] perform one-time and length-thresholded KV cache compression respectively, but do not distinguish between prefill and decoding. SnapStream builds on training-free compression literature by combining the one-time compression of SnapKV with efficient decoding using a fixed budget from StreamingLLM, fitting each technique into the prefill-decode scheduling of a production LLM server. We leave incorporating pre-computed, offloaded KV blocks/pages in the style of Quest and InfLLM to future work.
@@
However, training-free approaches invariably lead to some accuracy degradation, leading to a resurgence in trainable sparse attention. Cartridges [13] propose compressing large corpuses into trainable KV caches with much smaller sequence lengths that may be optionally loaded using cached prefix lookups. Titans [4] propose to augment Transformers with a neural memory module that aims to memorize the data seen during training and improve efficiency and accuracy on long context tasks. DeepSeek-V2 introduced Multi-Head Latent Attention (MLA) [10], which introduces shared low-rank projections of the query, key, and value vectors to reduce KV cache memory along the hidden dimension. However, doing so requires increased compute to project these compressed KVs back into the full hidden dimension during decoding. Native Sparse Attention (NSA) [35] proposes a similar KV cache structure to SnapStream, with hierarchical compression of past tokens, that is trainable and aims to reduce complexity for both training and inference. DeepSeek Sparse Attention (DSA) [12] proposes to finetune an existing dense attention model instead of training from scratch by training a token selection module, which still computes $O(L^{2})$ scores between tokens, but with fewer heads and smaller hidden dimensions than full attention, thus reducing the compute costs of prefill. While SnapStream already accommodates MLA by compressing the smaller, low-rank latent vectors along the sequence dimension instead of the full-rank KVs, admitting the other architectural modifications mentioned above requires additional modifications to fused kernels that we leave to future work.
VII Conclusion
@@
In this paper, we present SnapStream, a training-free KV cache compression method that employs SnapKV compression during prefill and a StreamingLLM rolling KV cache window during decoding. We describe the details of a static graph implementation of this method, as well as the details for a production 16-socket mixed tensor-parallel and data-parallel mapping of DeepSeek-R1-0528 on SN40Ls. This mapping uses SnapStream to reduce memory pressure and increase decoding throughput. We verify that SnapStream introduces minimal accuracy degradation on long sequence tasks for both Llama 3.1 8B Instruct and DeepSeek-R1, as well as additional reasoning and coding tasks for the latter model. Finally, we show how our SnapStream mapping enables us to achieve a $4.3\times$ improvement in decoding throughput, with at most a 5% increase in prefill latency. We hope this work inspires further study of KV cache compression methods in production LLM serving workloads and a variety of hardware accelerators, and leave a closer examination of hyperparameters and mappings of other sparse attention mechanisms to future work.